OpenStudy (anonymous):

MEDALS FOR ALLL ANSWERS!!!!!!!!!!!!!!!!!!!!! 4/5y+2=6

OpenStudy (anonymous):

4/(5y + 2) = 6/1 6(5y + 2) = 4 30y + 12 = 4 30y = -8 y = -8/30 and/ or (4/5)y + 2 = 6 (4/5)y = 4 y = 5